OpenStudy (anonymous):

show that: log (base a)x/log (base ab)x=1+log (base a)b

myininaya (myininaya):

did you try using change of base formula?

OpenStudy (anonymous):

ya..i did but somehow its nt comng

myininaya (myininaya):

\[\log_a(x)=\frac{\ln(x)}{\ln(a)}, \log_{ab}(x)=\frac{\ln(x)}{\ln(ab)}\]

myininaya (myininaya):

\[\frac{\ln(x)}{\ln(a)} \div \frac{\ln(x)}{\ln(ab)}=\frac{\ln(x)}{\ln(a)} \cdot \frac{\ln(ab)}{\ln(x)}\]

myininaya (myininaya):

now ask yourself does something cancel?

myininaya (myininaya):

and then also can you rewrite ln(ab)

OpenStudy (anonymous):

my subscription

myininaya (myininaya):

are you still stuck?

OpenStudy (anonymous):

yA PLEASE if posible mak it a litle clear...

OpenStudy (anonymous):

\[\text{Let } k =log_{ab}(x)\]\[\implies ab^k = x\]\[\implies log_a\left((ab)^k\right) = log_a(x)\]\[\implies log_a(a^k) + log_a(b^k) = log_a(x)\]\[\implies k + k(log_a(b)) = log_a(x)\]\[\implies \frac{log_a(x)}{k} = 1 + log_a(b)\] Substituting back in for k... \[\implies \frac{log_a(x)}{log_{ab}(x)} = 1 + log_a(b)\]

OpenStudy (anonymous):

Fun with logs!

OpenStudy (anonymous):

Err that first line should be \[(ab)^k\]

OpenStudy (anonymous):

Or the second.. whatever.

OpenStudy (anonymous):

realllllllllyyyyyyy it helped me

OpenStudy (anonymous):

thnk u

OpenStudy (anonymous):

Just be sure you understand what was done at each step.

OpenStudy (anonymous):

If not, please ask.

OpenStudy (anonymous):

ya ...i got it..

OpenStudy (anonymous):

\[\log_{a} x \log_{b} y \log_{c} z =\log_{b} x \log_{c} y \log_{a} z\]

OpenStudy (anonymous):

prove

OpenStudy (anonymous):

myininaya there is a factor by grouping problem waiting for you...